OpenStudy (anonymous):

If the current value of a property is $20,000 and its value is increasing at 8% per year. Find the expression for the value after t years. A. $20,000(1.08)t B. $20,000(1.8)t C. $20,000(0.92)t D. $20,000(1.18)t

OpenStudy (anonymous):

a

jimthompson5910 (jim_thompson5910):

\[\Large FV = P(1+r)^{t}\] \[\Large FV = 20000(1+0.08)^{t}\] \[\Large FV = 20000(1.08)^{t}\] So it looks like it's choice A